An electric flux of magnitude 180 passes through a flat horizontal surface that has an area 0.5 m2. If the electric field was uniform, what is the magnitude of the field if it passes the area at angle of 30° ?


1
Expert's answer
2021-06-21T11:42:24-0400

"\Phi=EA\cos\alpha\to E=\frac{\Phi}{A\cos\alpha}=\frac{180}{0.5\cdot \cos(90°-30°)}=720\ (V/m)" . Answer